please help.write the equation of the line


line y(2,11) line z(-1,-1) hope you understand this.Thanks for the help.

Answer by stanbon(75887)

write the equation of the line

line y(2,11) line z(-1,-1)
---------------------------
slope = (11--1)/(2--1) = 12/3 = 4
------------------
Form: y = mx + b and you know y=11 when x = 2, and m = 4; solve for b
--------------------
11 = 4*2 + b
b = 3
--------------------
EQUATION:
y = 4x + 3
==================
Cheers,
Stan H.